OpenStudy (anonymous):

express log 64/27 in terms of log2 and log3

OpenStudy (anonymous):

2 x + 7 < 7 x + 2 < 14 x - 1

OpenStudy (jdoe0001):

\(\huge ?!\)

OpenStudy (tkhunny):

\(64 = 2^{What}\)? \(27 = 3^{What}\)?

OpenStudy (jdoe0001):

notice -> http://www.chilimath.com/algebra/advanced/log/images/rules%20of%20exponents.gif <-- rule 2 listed there keep in mind that \(\bf 64=2^6\qquad \qquad 27=3^3\)